# Interpolating many-body wave functions for accelerated molecular dynamics on near-exact electronic surfaces

### v1

@@ Abstract
While there have been many developments in computational probes of both strongly-correlated molecular systems and machine-learning accelerated molecular dynamics, there remains a significant gap in capabilities between them, where it is necessary to describe the accurate electronic structure over timescales in which atoms move. We describe a practical approach to bridge these fields by interpolating the correlated many-electron state through chemical space, whilst avoiding the exponential complexity of these underlying states. With a small number of accurate correlated wave function calculations as a training set, we demonstrate provable convergence to near-exact potential energy surfaces for subsequent dynamics with propagation of a valid many-body wave function and inference of its variational energy at all points, whilst retaining a mean-field computational scaling. This represents a profoundly different paradigm to the direct interpolation of properties through chemical space in established machine-learning approaches. It benefits from access to all electronic properties of interest from the same model without relying on local descriptors, and demonstrates improved performance compared to the direct training on energies themselves. We combine this with modern systematically-improvable electronic structure methods to resolve the molecular dynamics for a number of correlated electron problems, including the proton dynamics of a Zundel cation trajectory, where we highlight the qualitative improvement from traditional machine learning or ab initio dynamics on mean-field surfaces.
Introduction
@@
The quantum fluctuations of interacting electrons represent the critical interaction between atoms which underpin all atomic bonding, dynamics and reactivity. Computational approaches for systems with strongly interacting electrons have undergone a number of major developments in recent decades, as emerging methods enable a description of correlated electronic structure for ever larger and more realistic systems with unprecedented accuracy [PhysRevX.7.031059, PhysRevX.10.011041, doi:10.1021/acs.jpclett.0c02621]. These modern approaches across both chemical and materials science include those based on tensor networks [10.1063/1.4905329, 10.1063/5.0050902, Larsson2022TheCD, doi:10.1126/science.abm2295, Li2019], stochastic methods [Booth2012TowardsAE, 10.1063/5.0150706, Ren2023, doi:10.1021/acs.jctc.1c01162], selected configuration interaction [KOCH1993193, 10.1063/1.5063376, 10.1063/1.1679199] and machine-learning inspired wave function ansatze [Choo_2020, https://doi.org/10.48550/arxiv.2109.12606, zhaoScalableNeuralQuantum2022, https://doi.org/10.48550/arxiv.2301.03755, Yang2020, PhysRevB.107.205119, bortone2023impact, PhysRevResearch.2.033429, Hermann2019DeepneuralnetworkSO, gerard2022goldstandard]. These approaches have allowed for the near-exact solution to the quantum many-body problem in these systems, providing high-accuracy insights for a few fixed atomic configurations, but have in general had little or no impact on our understanding of the physics and chemistry on the timescales of atomic and molecular motion.
@@
The reasons for this are obvious; while a small number of single point calculations with fixed atoms are possible, the different timescales of atomic dynamics and electronic quantum fluctuations mean that on the order of at least thousands of sequential electronic structure calculations are required. This is essential to propagate the atoms in molecular systems to relevant timescales, entailing generally prohibitive computation expense for these high-accuracy methods. This is particularly challenging for these emerging methods which can lack a ‘black-box’ use, requiring care to ensure reliable convergence at each point, while often also lacking analytic atomic forces to propagate the nuclear coordinates in time [doi:10.1021/acs.chemrev.9b00496]. For this reason ‘ab-initio Born-Oppenheimer molecular dynamics’ (AI-BOMD), where the atoms are classically propagated according to the potential energy surface of the electrons, is almost synonymous with a more empirical density functional description of the electronic structure (with exceptions [doi:10.1021/jz401931f, 10.3389/fmats.2015.00029, 10.1063/1.4941091]), which lacks systematic improvability and has many well documented deficiencies [marx2000ab]. These include an over-stabilisation of delocalized electronic states, as well as often inaccurate descriptions of dispersion forces, transition states or bond breaking among others [Cohen2012ChallengesFD]. These are critical parts of the phase space in real chemical dynamics, and the acute need for more reliable potential energy surfaces which build on the developments in accurate electronic structure is clear.
@@
The most widespread and successful resolution to this need has come from the ‘machine learning’ of force fields [Kabylda2023EfficientID, Fu2022ForcesAN, Unke2020MachineLF, Gkeka2020MachineLF]. These interpolate across chemical space between accurate single-point estimates of the electronic energy, based on local descriptors of the environment of each atom [doi:10.1021/acs.chemrev.1c00021]. While this approach to straddling the electronic and atomic timescales has been arguably one of the most successful contributions of machine learning to quantum-level simulations to date, it is not without its own drawbacks [Kulik_2022]. In particular, the local nature of the descriptors can lead to difficulty describing long-range interactions [doi:10.1021/acs.jctc.3c00704], as well as ‘holes’ where non-variational inferred energy estimates can lead to a collapse in the statistical sampling of phase space to these unphysical minima [doi:10.1021/acs.jctc.1c00647]. On a more fundamental level, since these approaches integrate out the electronic structure, there is in general no fundamental electronic variable at each sampled point (such as the wave function), meaning that the electronic properties which can be extracted are limited to the ones which correspond to the model definition. If the evolution of e.g. the dipole moment or charge distribution across a trajectory was desired this would not be accessible from a force field, and extensions to non-adiabatic effects are also far from straightforward in this framework, noting however significant recent research in these directions [Schtt2019UnifyingML, Fedik2022ExtendingML, doi:10.1021/acs.jpclett.0c00527].
@@
We take a different perspective and show that rather than interpolating observables such as the potential energy, we can instead interpolate the many-body electronic wave function itself through chemical space. Importantly, despite the many-body wave function of each training point being in general exponentially complex, inference of properties from the model can be achieved in a scaling which is the same as (hybrid) density functional theory, rendering this a practical scheme. This decouples the unfavourable scaling of high accuracy single point electronic structure calculations from the evaluation of the interpolated potential energy surface, and thus allows for the use of these electronic structure methods for atomic dynamics on realistic timescales. We show that the resulting potential energy surfaces for molecular dynamics over relevant timescales are systematically improvable to near-exactness via interpolation between highly accurate training configurations. Since a valid correlated many-electron state is propagated through chemical space, this paradigm enables all electronic properties of interest to be simultaneously accessible within the same model, without relying on local or low-rank descriptors. Furthermore, since the energy is computed as a rigorous quantum expectation value over this inferred state it provides a fully variational potential energy estimate (precluding ‘holes’) for all atomic configurations, allows for clear evidence of systematic improvability to exactness as the training set is enlarged, an inductive bias of the model away from poorly described regions of phase space, and simple access to analytic atomic forces of the model for efficient propagation of dynamics.
@@
We combine this approach for interpolating wave functions with modern density matrix renormalization group (DMRG) methods, allowing convergence of the strongly correlated potential energy surfaces to near exactness [10.1063/1.4905329, 10.1063/5.0050902]. We demonstrate that this can provide a fully correlated electronic description of reactive molecular dynamics beyond traditional parameterized or machine-learned force fields, over time scales which would be inaccessible without this acceleration. We show this can result in qualitative differences in behavior for a number of proto-typical molecular dynamics simulations compared to both density functional and traditional machine-learned force field approaches [10.1063/5.0160898]. Finally, we consider a high-energy reaction trajectory on a near-exact potential surface for the Zundel cation, a key intermediate for the Grotthuss mechanism for hydrogen diffusion through aqueous solutions [Agmon1995TheGM, Cukierman2006EtTG, 10.1063/1.4941091]. We find a hydrogen atom ejected from the system over the order of a hundred femtoseconds, contradicting widely employed density function results, and demonstrating a marked difference in the charge distribution over the ions.
@@
where $C_{n_{1},n_{2},\dots,n_{L}}$ is the rank-$L$ tensor of probability amplitudes over the electronic configurations, and $n_{i}$ indexes the four local Fock states of the $i^{\textrm{th}}$ orbital; either unoccupied, spin-up, spin-down, or doubly occupied with electrons for each orbital. In general, both the probability amplitudes, and the single-particle orbitals defining each electronic configuration $|n_{1},n_{2},\dots,n_{L}\rangle$ will change with atomic configuration ${\bf R}$. However, we aim to represent an approximation to the correlated electronic state at a different chemical configuration (and therefore electronic Hilbert space) with the same tensor of probability amplitudes over these electronic states. We exploit the fact that the properties of the exact state are invariant to orthogonal rotations of the single-particle orbitals, but that the probability amplitudes themselves will vary with this choice. Therefore, to enable transferrability between chemical environments, we seek a choice of orbital representation in which the probability amplitudes of the exact many-electron state change least between atomic configurations of interest.
@@
where $\mathbf{n}$ denotes the many-electron configurations in the SAO basis of the test geometry, with $|\mathbf{n}\rangle\equiv|n_{1},n_{2},\dots,n_{L}\rangle$, and with $C_{\mathbf{n}}^{(a)}$ and $C_{\mathbf{n}}^{(b)}$ the fixed SAO probability amplitudes of the training points at atomic geometries $a$ and $b$ respectively. This definition therefore implicitly transfers the probability amplitudes between the Hilbert spaces of the training and test geometries. The $K_{ijkl}(\mathbf{R})$ tensor is the two-electron reduced Hamiltonian defined in the SAO basis of the test geometry (given explicitly in the Methods section) with second-quantized Fermionic operators acting in this basis as ${\hat{c}}^{(\dagger)}$ shown.
Since the Hamiltonian is a sum of only two-electron interactions, the contraction over the exponential many-electron configurations, $\mathbf{n}$, is performed for all pairs of training states to give the transition two-body density matrices $\boldsymbol{\Gamma}^{ijkl}$ defined in Eq. 3. Crucially, since the training probability amplitudes are defined not to change with test geometry, this contraction is only performed once on the training states, not for each test geometry. The construction of the subspace Hamiltonian at a test geometry therefore only requires the $\mathcal{O}[L^{4}]$ contraction of Eq. 3 – the same scaling as traditional (hybrid) density functional theory, with only $K_{ijkl}(\mathbf{R})$ changing with test configuration. This scaling for evaluation of the model at test geometries could also be further lowered with factorizations exploiting the low-rank nature of the tensors in Eq. (3[doi:10.1021/acs.jctc.7b00605, 10.1063/1.5047207, doi:10.1021/acs.jctc.8b00780, doi:10.1021/acs.jctc.3c00851]. The overlap between the training states, $\boldsymbol{\mathcal{S}}$ of Eq. 2, can similarly be precomputed during the training phase, as
@@
due to the orthonormality of the SAO basis at all test geometries. This ensures that the overlaps of the training states do not change with geometry, despite the physical training wave functions changing with atomic rearrangements as they transform between Hilbert spaces. In this way, the exponential complexity of the many-electron states are completely avoided in the inference of wave functions at new test points, by representing the training states in the polynomially-compact tensors $\boldsymbol{\Gamma}^{ijkl}$ and their overlaps.
@@
The lowest energy eigenvector from the diagonalization of this Hamiltonian (whose dimensionality scales only with the number of training points and is independent of system size) defines the specific variationally optimal linear combination of training probability amplitudes for the state, which can subsequently be used to predict any electronic property at this test geometry. Due to the variationality we have the desirable properties that the inferred state at a point which coincides with a training geometry must necessarily be exact, as well as the fact that each additional training point must necessarily lower the inferred energy towards the exact electronic solution across all possible test atomic configurations, assuming linear independence. In this way, the method more closely resembles a reduced order method than a machine learning model, yet both are useful viewpoints. Due to the fixed training amplitudes across geometries, as well as the variational optimization of the model, computing analytic atomic forces from the inferred state is also straightforward via the Hellmann-Feynman theorem (see Methods), ensuring a particular relevance of this acceleration in molecular dynamics applications.
@@
Figure 1: Schematic overview of the proposed ‘Eigenvector Continuation’ scheme for the high accuracy prediction across chemical space from few ab initio data points. Utilizing training states in a Löwdin atomic orbital (SAO) basis, states are predicted for new molecular geometries by diagonalization in the subspace of the training states, allowing for a fast variational prediction of highly accurate potential energy surfaces and force fields via the compact one- and two-body transition density matrices between states. The scheme is exemplified for symmetrically stretched chains of $6$ hydrogen atoms in a minimal basis set (STO-6G) using one, two, and three training points. $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\Gamma}^{ijkl}$ represent the geometry-agnostic overlap and two-body transition density matrices between training states. $\langle\mathbf{n}|\Psi^{(a)}\rangle$ shows the value of the (exponentially many) wavefunction amplitudes in the SAO basis for the $a^{\mathrm{th}}$ training state, and the test geometry shows the combination of training states used to define its own SAO probability amplitudes.
@@
This approach also builds on the perspective of ‘eigenvector continuation’ which was recently introduced in both nuclear physics and condensed matter lattice models, where an eigenstate is analytically continued to different parts of the phase diagram [PhysRevLett.121.032501, PhysRevC.101.041302, duguet2023eigenvector, KONIG2020135814, PhysRevC.107.064316, DRISCHLER2021136777, 10.1063/5.0141145]. Even more recently this was extended to simple ab initio quantum chemistry applications, with a related scheme to the one proposed [mejutozaera2023quantum]. However, a crucial difference was the use of a non-orthogonal atomic basis, which necessitated evaluation of the test point Hamiltonian directly from the many-body states. This retained the exponential complexity of the many-body state for inference at each test geometry, which is avoided here. The method of Mejuto-Zaera et. al. was therefore presented instead as an approach for quantum computers, where many-body unitary operations can be applied in polynomial complexity. In contrast, the SAO basis for the interpolation formally breaks this requirement, ensuring the approach is amenable to classical computation in the predictions at test points with tractable mean-field computational cost.
@@
This approach is invariant to translation and rigid body rotations of the chemical system, provided a consistent ordering of the SAO representation is maintained, which is straightforward to achieve. However, this is not trivial for atomic permutations or point group symmetries which would change the SAO ordering, and hence probability amplitudes of the state definition. Furthermore, in contrast with building a force field based on local descriptors, the inference requires the same dimensionality Hilbert space for the electronic state, necessitating that the training and prediction points are taken from the same sized system. Future work will look to relax this constraint.
@@
While the proof-of-principle in Fig. 1 demonstrates excellent accuracy with few training points, it is also interpolation within a simple one-dimensional phase space of geometries. We now compare to a far larger phase space, composed of averaging the errors in both the inferred energy and analytic forces on the atoms over randomly oriented three-dimensional displacements of each atom from a ten-atom linear hydrogen chain. This provides an exponentially large phase space of distorted chain configurations to test, where the radius of the displacements of each atom can be used to control the magnitude of the geometric distortions from the parent linear chain. Only five training points from the symmetric stretch of the equidistant linear chain are used. We consider the increase in error as the magnitude of the displacements are increased in Fig. 2, as the test configurations move further from these training samples. We also compare these errors to a Gaussian approximated potential (GAP); a widely used machine-learning approach based on Gaussian process regression in a space of local descriptors from the superposition of atomic potentials [PhysRevLett.104.136403, 10.1063/5.0160898]. This models a force field directly from the same training energies, but results in a materially larger error for the energy and forces over all displacements. We note that five points would generally be a very small training set for GAP, and that improved techniques to directly train on the forces of the training data themselves or improved model definitions were not used [PhysRevB.95.214302]. Nevertheless, a demonstration that inferring the wave function amplitudes themselves can outperform traditional machine-learning inference of the properties directly is noteworthy. Furthermore, we compare to Hartree–Fock theory (HF), which neglects all correlated electron effects and has the same computational scaling as the inference of the proposed ‘eigenvector continuation’ scheme. This is also significantly worse at small distortions of the chain, though outperforms the largest distortions which are far from the training geometries and deep in the extrapolation regime.
@@
We consider these approaches for constructing training states and the subsequent MD of a water molecule in increasing basis sets in Fig. 3. In particular, we consider convergence of the predicted vibrational frequency of the $a_{1}$ symmetric stretching mode as the number of training points increases. For the smallest basis, we find the full vibrational dynamics converge with just three training points, where we can compare directly to exact FCI calculations of the dynamics. As we increase the basis, FCI is intractable and we restrict the training to a CAS of low-energy orbitals, where the number of training points required grows modestly to seven and thirteen training points in a cc-pVDZ and cc-pVTZ basis respectively. While we compare to CAS-CI for these larger bases for a reference estimate, we note that the inferred energies are necessarily variationally below these approximate training states, as shown in the extended data.

### v2

@@ Abstract
While there have been many developments in computational probes of both strongly-correlated molecular systems and machine-learning accelerated molecular dynamics, there remains a significant gap in capabilities in simulating accurate non-local electronic structure over timescales on which atoms move. We develop an approach to bridge these fields with a practical interpolation scheme for the correlated many-electron state through the space of atomic configurations, whilst avoiding the exponential complexity of these underlying electronic states. With a small number of accurate correlated wave functions as a training set, we demonstrate provable convergence to near-exact potential energy surfaces for subsequent dynamics with propagation of a valid many-body wave function and inference of its variational energy whilst retaining a mean-field computational scaling. This represents a profoundly different paradigm to the direct interpolation of potential energy surfaces in established machine-learning approaches. We combine this with modern electronic structure approaches to systematically resolve molecular dynamics trajectories and converge thermodynamic quantities with a high-throughput of several million interpolated wave functions with explicit validation of their accuracy from only a few numerically exact quantum chemical calculations. We also highlight the comparison to traditional machine-learned potentials or dynamics on mean-field surfaces.
@@
The quantum fluctuations of interacting electrons represent the critical interaction between atoms which underpin all atomic bonding, dynamics and reactivity. Computational approaches for systems with strongly interacting electrons have undergone a number of major developments in recent decades, as emerging methods enable a description of correlated electronic structure for ever larger and more realistic systems with unprecedented accuracy [PhysRevX.7.031059, PhysRevX.10.011041, doi:10.1021/acs.jpclett.0c02621]. These modern approaches across both chemical and materials science include those based on tensor networks [10.1063/1.4905329, 10.1063/5.0050902, Larsson2022TheCD, doi:10.1126/science.abm2295, Li2019], stochastic methods [Booth2012TowardsAE, 10.1063/5.0150706, Ren2023, doi:10.1021/acs.jctc.1c01162], selected configuration interaction [KOCH1993193, 10.1063/1.5063376, 10.1063/1.1679199] and machine-learning inspired wave function ansatze [Choo_2020, https://doi.org/10.48550/arxiv.2109.12606, zhaoScalableNeuralQuantum2022, https://doi.org/10.48550/arxiv.2301.03755, Yang2020, PhysRevB.107.205119, bortone2023impact, PhysRevResearch.2.033429, Hermann2019DeepneuralnetworkSO, gerard2022goldstandard]. This has allowed for the near-exact solution to the quantum many-electron problem in these systems, providing high-accuracy insights for a few fixed atomic configurations, but have in general had little or no impact on our understanding of the physics and chemistry on the timescales of atomic and molecular motion.
@@
The reasons for this are obvious; while a small number of single point calculations with fixed nuclei are possible, the different timescales of atomic dynamics and electronic quantum fluctuations mean that on the order of at least thousands of sequential electronic structure calculations are required. This is essential to propagate the atoms in molecular systems to relevant timescales, entailing generally prohibitive computation expense for these high-accuracy methods. This is particularly challenging for these emerging methods which can lack a ‘black-box’ use, requiring care to ensure reliable convergence at each point, while often also lacking analytic atomic forces to propagate the nuclear coordinates in time [doi:10.1021/acs.chemrev.9b00496]. Important developments have been made in recent years in extending the application of established ground-state quantum chemical models to atomic dynamics [doi:10.1021/jz401931f, 10.3389/fmats.2015.00029, 10.1063/1.4941091, C4CP05192K, Hutter2018, 10.1063/5.0212274], while ‘active space’ methods are also increasingly widely used for stronger correlation or excited state molecular dynamics [doi:10.1021/acs.chemrev.7b00577]. However, the additional cost of these approaches have meant that ‘ab-initio Born-Oppenheimer molecular dynamics’ (AI-BOMD), where the atoms are classically propagated according to the potential energy surface of the electrons, is almost synonymous with a more empirical density functional description of the electronic structurewhich lacks systematic improvability and has many well documented deficiencies [PhysRevLett.55.2471, doi:10.1073/pnas.0500193102, marx2000ab]. These include an over-stabilisation of delocalized electronic states, as well as often inaccurate descriptions of dispersion forces, transition states or bond breaking among others [Cohen2012ChallengesFD]. These are critical parts of the phase space in real chemical dynamics, and the acute need for more reliable potential energy surfaces which build on the developments in accurate electronic structure is clear.
@@
The most widespread and successful resolution to this need has come from the ‘machine learning’ of force fields [Kabylda2023EfficientID, Fu2022ForcesAN, Unke2020MachineLF, Gkeka2020MachineLF]. These interpolate across chemical space between accurate single-point estimates of the electronic energy, based on local descriptors of the environment of each atom [doi:10.1021/acs.chemrev.1c00021]. While this approach to straddling the electronic and atomic timescales has been arguably one of the most successful contributions of machine learning to quantum-level simulations to date, it is not without its own drawbacks [Kulik_2022]. In particular, the local nature of the descriptors can lead to difficulty describing long-range interactions [doi:10.1021/acs.jctc.3c00704], as well as ‘holes’ where non-variational inferred energy estimates can lead to a collapse in the statistical sampling of phase space to these unphysical minima [doi:10.1021/acs.jctc.1c00647]. On a more fundamental level, since these approaches integrate out the electronic structure, there is in general no fundamental electronic variable at each sampled point (such as the wave function), meaning that the electronic properties which can be extracted are limited to the ones which correspond to the model definition. If the evolution of e.g. the dipole moment or charge distribution across a trajectory was desired this would not be accessible from a force field, and extensions to non-adiabatic effects are also far from straightforward in this framework, noting however significant recent research in these directions [Schtt2019UnifyingML, Fedik2022ExtendingML, doi:10.1021/acs.jpclett.0c00527, D3FD00113J].
@@
We take a different perspective and show that rather than interpolating observables such as the potential energy, we can instead interpolate the many-body electronic wave function itself through the phase space of molecular conformations. Importantly, despite the many-body wave function of each training point being in general exponentially complex, inference of properties from the model can be achieved in a scaling which is the same as (hybrid) density functional theory, rendering this a practical scheme. This decouples the unfavourable scaling of high accuracy single point electronic structure calculations from the evaluation of the interpolated potential energy surface, and thus allows for the use of these electronic structure methods for molecular dynamics on realistic timescales. We show that the resulting potential energy surfaces and molecular dynamics are systematically improvable to near-exactness via interpolation between highly accurate training configurations. Since a valid correlated many-electron state is propagated through the sampled phase space, this paradigm enables all electronic properties of interest to be simultaneously accessible within the same model, without relying on local or low-rank descriptors. Furthermore, since the energy is computed as a rigorous quantum expectation value over this inferred state, it provides a fully variational potential energy estimate (precluding ‘holes’) for all atomic configurations, allows for clear evidence of systematic improvability to exactness as the training set is enlarged, an inductive bias of the model away from poorly described regions of phase space, and simple access to analytic atomic forces of the model for efficient propagation of dynamics.
@@
We combine this approach for interpolating wave functions with modern density matrix renormalization group (DMRG) methods, allowing convergence of the strongly correlated potential energy surfaces to near exactness within the employed basis [10.1063/1.4905329, 10.1063/5.0050902]. We demonstrate that this can provide a fully correlated electronic description of reactive molecular dynamics beyond traditional parameterized or machine-learned force fields, and ensembles of thermalized trajectories for equilibrated quantities over time scales which would be inaccessible without this acceleration. We show this can result in qualitative differences in behavior for a number of proto-typical molecular dynamics simulations compared to both density functional and traditional machine-learned force field approaches [10.1063/5.0160898]. Finally, we compute both thermalized expectation values from canonical emsembles and reactive high-energy dynamical trajectories on a near-exact potential surface for the Zundel cation, a key intermediate for the Grotthuss mechanism for hydrogen diffusion through aqueous solutions [Agmon1995TheGM, Cukierman2006EtTG, 10.1063/1.4941091]. With explicit validation of the accuracy of the surface, we compare the dynamics to both density functional theory results and other quantum chemical methods for both for structural and electronic quantities, highlighting marked differences which can result from the quality of the surface.
Results and discussion
@@
where $C_{n_{1},n_{2},\dots,n_{L}}$ is the rank-$L$ tensor of probability amplitudes over the electronic configurations, and $n_{i}$ indexes the four local Fock states of the $i^{\textrm{th}}$ orbital; either unoccupied, spin-up, spin-down, or doubly occupied with electrons for each orbital. In general, both the probability amplitudes, and the single-particle orbitals defining each electronic configuration $|n_{1},n_{2},\dots,n_{L}\rangle$ will change with atomic configuration ${\bf R}$. However, we aim to represent an approximation to the correlated electronic state at a different atomic configuration (and therefore electronic Hilbert space) with the same tensor of probability amplitudes over these electronic states. We exploit the fact that the properties of the exact state are invariant to orthogonal rotations of the single-particle orbitals, but that the probability amplitudes themselves will vary with this choice. Therefore, to enable transferrability between chemical environments, we seek a choice of orbital representation in which the probability amplitudes of the exact many-electron state change least between atomic configurations of interest.
@@
where $\mathbf{n}$ denotes the many-electron configurations in the SAO basis of the test geometry, with $|\mathbf{n}\rangle\equiv|n_{1},n_{2},\dots,n_{L}\rangle$, and with $C_{\mathbf{n}}^{(a)}$ and $C_{\mathbf{n}^{\prime}}^{(b)}$ the fixed SAO probability amplitudes of the training wave functions at atomic geometries $a$ and $b$ respectively. This definition therefore implicitly transfers the probability amplitudes between the Hilbert spaces of the training and test geometries. The $K_{ijkl}(\mathbf{R})$ tensor is the two-electron reduced Hamiltonian defined in the SAO basis of the test geometry (given explicitly in the Methods section) with second-quantized Fermionic operators acting in this basis as ${\hat{c}}^{(\dagger)}$ shown.
Since the Hamiltonian is a sum of only two-electron interactions, the contraction over the exponential many-electron configurations, $\mathbf{n}$, is performed for all pairs of training states to give the transition two-body density matrices $\boldsymbol{\Gamma}^{ijkl}$ defined in Eq. 3. Crucially, since the training probability amplitudes are defined not to change with test geometry, this contraction is only performed once on the training states, not for each test geometry. The construction of the subspace Hamiltonian at a test geometry therefore only requires the $\mathcal{O}[L^{4}]$ contraction of Eq. 3, with only the $K_{ijkl}(\mathbf{R})$ term changing with test configuration. The overlap between the training states, $\boldsymbol{\mathcal{S}}$ of Eq. 2, can similarly be precomputed during the training phase, as
@@
due to the orthonormality of the SAO basis at all test geometries. This ensures that the overlaps of the training states do not change with geometry, despite the physical training wave functions changing with atomic rearrangements as they transform between Hilbert spaces. In this way, the exponential complexity of the many-electron states are completely avoided in the inference of wave functions at new test points, by representing the training states in the polynomially-compact tensors $\boldsymbol{\Gamma}^{ijkl}$ and their overlaps. The inference of the model requires a computational scaling of non-iterative $\mathcal{O}[N^{2}L^{4}]$ after the density matrices of the training states have been precomputed in the training stage – the same formal scaling with system size as traditional (hybrid) density functional theory. This scaling for evaluation of the model at test geometries could also be further lowered with factorizations exploiting the low-rank nature of the $\Gamma_{ab}^{ijkl}$ tensors [doi:10.1021/acs.jctc.7b00605, 10.1063/1.5047207, doi:10.1021/acs.jctc.8b00780, doi:10.1021/acs.jctc.3c00851].
@@
The lowest energy eigenvector from the diagonalization of this Hamiltonian (whose dimensionality scales only with the number of training points and is independent of system size) defines the specific variationally optimal linear combination of training probability amplitudes for the state, which can subsequently be used to predict any electronic property at this test geometry. Due to the variationality we have the desirable properties that the inferred state at a point which coincides with a training geometry must necessarily be exact, as well as the fact that each additional training point must necessarily lower the inferred energy towards the exact electronic solution across all possible test atomic configurations, assuming linear independence. In this way, the method more closely resembles a reduced order method than a machine learning model, where we define the Hamiltonian in a subspace defined by a fixed set of many-body vectors taken from training wave functions at different geometries, yet both are useful viewpoints. Due to the fixed training amplitudes across geometries, as well as the variational optimization of the model, computing analytic atomic forces from the inferred state is also straightforward (see Methods), ensuring a particular relevance of this acceleration in molecular dynamics applications.
@@
Figure 1: Schematic overview of the proposed ‘Eigenvector Continuation’ scheme for the high accuracy prediction across conformational space from few ab initio data points. Utilizing training states in a Löwdin atomic orbital (SAO) basis, states are predicted for new molecular geometries by diagonalization in the subspace of the training states, allowing for a fast variational prediction of highly accurate potential energy surfaces and force fields via the compact one- and two-body transition density matrices between states. The scheme is exemplified for symmetrically stretched chains of $6$ hydrogen atoms in a minimal basis set (STO-6G) using one, two, and three training points. $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\Gamma}^{ijkl}$ represent the geometry-agnostic overlap and two-body transition density matrices between training states. $\langle\mathbf{n}|\Psi^{(a)}\rangle$ shows the value of the (exponentially many) wavefunction amplitudes in the SAO basis for the $a^{\mathrm{th}}$ training state, and the test geometry shows the combination of training states used to define its own SAO probability amplitudes.
@@
This approach also builds on the perspective of ‘eigenvector continuation’ which was recently introduced in both nuclear physics and condensed matter lattice models, where an eigenstate is analytically continued to different parts of the phase diagram [PhysRevLett.121.032501, PhysRevC.101.041302, duguet2023eigenvector, KONIG2020135814, PhysRevC.107.064316, DRISCHLER2021136777, 10.1063/5.0141145, D4FD00062E]. Even more recently this was extended to simple ab initio quantum chemistry applications, with a related scheme to the one proposed [mejutozaera2023quantum]. However, a crucial difference was the use of a non-orthogonal atomic basis, which necessitated evaluation of the test point Hamiltonian directly from the many-body states. This retained the exponential complexity of the many-body state for inference at each test geometry, which is avoided here. The method of Mejuto-Zaera et. al. was therefore presented instead as an approach for quantum computers, where many-body unitary operations can be applied in polynomial complexity. In contrast, the SAO basis for the interpolation formally breaks this requirement, ensuring the approach is amenable to classical computation in the predictions at test points with tractable mean-field computational cost.
@@
This approach is invariant to translation and rigid body rotations of the chemical system, provided a consistent ordering of the SAO representation is maintained, which is straightforward to achieve. However, this is not trivial for atomic permutations or point group symmetries which would change the SAO ordering, and hence probability amplitudes of the state definition. Furthermore, in contrast with building a force field based on local descriptors, the inference requires the same dimensionality Hilbert space for the electronic state, necessitating that the training and prediction points are taken from the same sized system. This is a significant difference to force field approaches with local representations which allow for scaling the system size after training, ensuring a different scope of applicability to the proposed approach [doi:10.1021/acs.chemrev.1c00021]. Future work will look to relax this constraint.
@@
While the proof-of-principle in Fig. 1 demonstrates excellent accuracy with few training points, it is also interpolation within a simple one-dimensional phase space of geometries. We now compare to a far larger phase space, composed of averaging the errors in both the inferred energy and analytic forces on the atoms over randomly oriented three-dimensional displacements of each atom from a ten-atom linear hydrogen chain. This provides an exponentially large phase space of distorted chain configurations to test, where the radius of the displacements of each atom can be used to control the magnitude of the geometric distortions from the parent linear chain from which the training data is obtained. Only five training points from the symmetric stretch of the equidistant linear chain are used. We consider the increase in error as the magnitude of the displacements are increased in Fig. 2, as the test configurations move further from these training samples. We also compare these errors to a Gaussian approximated potential (GAP); a widely used machine-learning approach based on Gaussian process regression in a space of local descriptors from the superposition of atomic potentials [PhysRevLett.104.136403, 10.1063/5.0160898]. This models a force field directly from the same training energies, but results in a materially larger error for the energy and forces over all displacements. We note that five points would generally be a very small training set for GAP, and that improved techniques to directly train on the forces of the training data themselves or improved model definitions were not used [PhysRevB.95.214302, Kozinsky21].
Nevertheless, a demonstration that inferring the wave function amplitudes themselves can outperform traditional machine-learning inference of the properties directly is noteworthy. Furthermore, we compare to Hartree–Fock theory (HF), which neglects all correlated electron effects and has the same computational scaling as the inference of the proposed ‘eigenvector continuation’ scheme. This is also significantly worse at small distortions of the chain, though outperforms the largest distortions which are far from the training geometries and deep in the extrapolation regime.
@@
We consider these approaches for constructing training states and the subsequent MD of a water molecule in increasing basis sets in Fig. 3. In particular, we consider convergence of the predicted vibrational frequency of the $a_{1}$ symmetric stretching mode as the number of training points increases. For the smallest basis, we find the full vibrational dynamics converge with just three training points, where we can compare directly to exact FCI calculations of the dynamics. As we increase the basis, FCI is intractable and we restrict the training to a CAS of low-energy orbitals, where the number of training points required grows modestly to seven and thirteen training points in a cc-pVDZ and cc-pVTZ basis respectively. While we compare to CASCI for these larger bases for a reference estimate, we note that the inferred energies are necessarily variationally below these approximate training states. Furthermore, any discontinuities in the potential energy which arise commonly due to abrupt changes in the in the active space character are necessarily removed in the interpolated surface. Both of these features are elaborated on in the extended data.